14. (a) A photon with initial wavelength of 6.76×10−12 m strikes an electron, and then has a wavelength of 7.02×10−12 m after it is deflected. What is the deflection angle?

(b) A photon with initial wavelength of 7.36×10−12 m strikes an electron. What is the maximum possible wavelength of the deflected photon and what angle does this correspond to?
Answer #1

According to Compton effect,

Δλ = λ/ - λ = (h/mec) (1 - cosθ)= 2.43 × 10-12×(1- cos)

As h/mec = 2.43 x 10-12 m

a) (7.02-6.76) × 10-12 = 2.43 × 10-12 (1-cos)

1-cos = 0.107

Cos = 0.893

= 26.750

b) maximum change in wavelength takes place for deflection angle of 1800

Δλ = λ/ - λ = (h/mec) (1 - cosθ)

For = 1800

Δλ = λ/ - λ = (h/mec) × 2 = 2× 2.43 x 10-12 m

λ/ - 7.36×10-12 m = 4.86 × 10-12

λ/ = 1.222 ×10-11 m